# Non-commutative Iwasawa theory of abelian varieties over global function fields

### v1

@@ 5. Akashi series
where the sum runs over places $w$ of $F^{\mathrm{cyc}}$ lying above $v$, and we also write $w$ for a fixed prime in $F_{\infty}$ lying above $w$. We claim that $J_{v}(A/F_{\infty})=0$ for $v\in S^{\prime}$. Indeed, if we let $w$ denote a prime of $F_{\infty}$ above $v$, then $\dim G_{w}=2$ by our assumption that $v\in S^{\prime}$. The theorem of Iwasawa [28, Theorem 7.5.3] then tells us that $F_{\infty,w}$ has no nontrivial $p$-extension. Hence $H^{1}(F_{\infty,w},A_{p^{\infty}})=0$ and this establishes our claim.
Taking this claim into account, it remains to compute $\mathrm{Ak}\left(J_{v}(A/F_{\infty})\right)$ for $v\in S\setminus S^{\prime}$. For a prime $v$ of $F$, we write $w$ for a prime of $F^{\mathrm{cyc}}$ lying above $v$, and by abuse of notation, we also write $w$ for a fixed prime in $F_{\infty}$ lying above $w$. We then denote by $H_{w}$ the decomposition group in $H$. Shapiro’s lemma gives
@@
If $v\in S\>\backslash\>S^{\prime}$, then $\mathrm{cd}_{p}(H_{w})=0$. Thus, the right hand side of (5.6) vanishes for all $i\geq 1$. For the $J_{v}(A/F_{\infty})^{H}$ term, one can check easily that the restriction map
@@
is an isomorphism using the fact that $\mathrm{cd}_{p}(H_{w})=0$. In conclusion, we have
$\mathrm{Ak}\Big{(}J_{v}(A/F_{\infty})\Big{)}=\left\{\begin{array}[]{ll}f_{J_{v}(A/F^{\mathrm{cyc}})}&\hbox{if $v\in S\setminus S^{\prime}$;}\\ 0,&\hbox{if $v\in S^{\prime}$.}\end{array}\right.$
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$. Let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then we have
@@
Putting the calculations of Lemma 5.3 and 5.4 into (5.3), we obtain
@@
Thanks to Theorem 4.2, there is a short exact sequence
@@
where all the terms are cofinitely generated over ${\mathbb{Z}}_{p}$ by Lemma 3.3 and Theorem 3.4. Consequently, we have
@@
Combining all of the above, we obtain the required conclusion of the theorem. ∎
We record an interesting corollary of our Theorem 5.5 which is an analogue of a result of Howson [15] (also see [6]).
@@
Retain the settings of Theorem 5.5. Suppose further that $G$ is pro-$p$ with dimension $\geq 2$. We then have the following formula
$\begin{split}\mathrm{rank}_{{\mathbb{Z}}_{p}[[H]]}X(A/F_{\infty})=\mathrm{rank}_{{\mathbb{Z}}_{p}}X(A/F^{\mathrm{cyc}})-\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}\\ +\sum_{w\in S^{\prime}(F^{\mathrm{cyc}})}\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w})^{\vee}.\end{split}$
@@
In the discussion of the proof, whenever $N$ is a finitely torsion ${\mathbb{Z}}_{p}[[\Gamma]]$-module with trivial $\mu$-invariant, we shall represent the characteristic series $f_{N}$ by a distinguished polynomial, and so it makes sense to speak of $\deg f_{N}$. In particular, we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(N)=\deg f_{N}$. Now, if $M$ is a ${\mathbb{Z}}_{p}[[G]]$-module which is finitely generated over ${\mathbb{Z}}_{p}[[H]]$, then $H_{i}(H,M)$ is finitely generated over ${\mathbb{Z}}_{p}$ for every $i$. Hence, we have
${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)=\sum_{i\geq 0}(-1)^{i}{\mathrm{rank}}_{{\mathbb{Z}}_{p}}H_{i}(H,M)$
(cf. [15, Theorem 1.1]). Combining this with the above, we see that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)$ can be read off from $\mathrm{Ak}(M)$. More precisely, we have
${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(M)=\sum_{i\geq 0}(-1)^{i}\deg f_{H_{i}(H,M)}.$
Now, rewriting the conclusion of Theorem 5.5, we have
@@
Since $\dim H\geq 1$ by hypothesis, we see that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(A_{p^{\infty}}(F_{\infty})^{\vee})=0$. It remains to show that
@@
for $v\in S^{\prime}$, where $w$ runs through the primes of $F^{\mathrm{cyc}}$ above $v$. For this, we note that $J_{v}(A/F^{\mathrm{cyc}})=\oplus_{w|v}H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})$. Since $v\in S^{\prime}$, it follows from the proof of Lemma 5.4 that $H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=H^{1}(H_{w},A_{p^{\infty}}(F_{\infty,w}))$. Since $\dim H_{w}\geq 1$ and ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H]]}(A_{p^{\infty}}(F_{\infty,w})^{\vee})=0$, we have
@@ 6. Control Theorem
In this section, we will prove the following control theorem which will allow us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group over ${\mathrm{Sel}}(A/F)$ over the base field $F$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic different from $p$. Let $F^{\mathrm{cyc}}$ denote the cyclotomic ${\mathbb{Z}}_{p}$-extension of $F$. Write $\Gamma_{n}={\mathrm{Gal}}(F_{n}/F)$, where $F_{n}$ for the $n$-th layer with of the ${\mathbb{Z}}_{p}$-extension $F^{\mathrm{cyc}}/F$ satisfying $[F_{n}:F]=p^{n}$. Then the restriction map
@@
has finite kernel and cokernel which are bounded independently of $n$.
@@
We will closely follow the method of the proof of the control theorem for fine Selmer groups over a number field as presented in [20]. We consider the following commutative diagram with exact rows
[…]
Since $\Gamma_{n}$ has $p$-cohomological dimension $1$, the inflation-restriction sequence tells us that $\ker\beta_{n}=H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))$ and $\mathrm{coker}~{}\!\beta_{n}=0$. Write $M=\left(A_{p^{\infty}}(F^{\mathrm{cyc}})\right)^{\vee}$. Then $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module with a continuous action by $\Gamma$, so that it is a finitely generated torsion $\Lambda(\Gamma_{n})$-module for every $n$. It follows by [28, Proposition 5.3.20] that
@@
But we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}M_{\Gamma_{n}}=0$ since $M_{\Gamma_{n}}=\left(A_{p^{\infty}}(F_{n})\right)^{\vee}$ is finite by the Mordell–Weil theorem. It follows that $M^{\Gamma_{n}}$ is also finite. Since $\Gamma_{n}$ has $p$-cohomological dimension $1$, we have
@@
which is the Pontryagin dual of $M^{\Gamma_{n}}$. Since $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module, the order of $M^{\Gamma_{n}}$ is bounded independently of $n$ by $M(p)$. It remains to show that $\ker\gamma_{n}$ is finite and bounded independently of $n$ in order to conclude that the same holds for $\ker\alpha_{n}$ and $\mathrm{coker}~{}\!\alpha_{n}$, by the snake lemma. For this, since $v\in S$ does not lie above $p$, it suffices to show that for each prime $w$ of $F^{\mathrm{cyc}}$ lying above $v\in S$, $H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))$ is finite and bounded independently of $n$. This can be shown by replacing $\Gamma_{n}$ with $\Gamma_{w}$ and $M$ with $M_{w}=\left(A_{p^{\infty}}(F^{\mathrm{cyc}}_{w})\right)^{\vee}$, which is still a finitely generated ${\mathbb{Z}}_{p}$-module, in the above argument showing that $H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))$ is finite and bounded independently of $n$. ∎
We have shown in Theorem 3.4 that $X(A/F^{\mathrm{cyc}})$ is a finitely generated torsion $\Lambda(\Gamma)$-module with $\mu$-invariant zero. Thus, the structure theorem for finitely generated $\Lambda(\Gamma)$-modules tells us that there is a pseudo-isomophism
@@
where we have identified $\Lambda(\Gamma)\simeq{\mathbb{Z}}_{p}[[T]]$ using our fixed topological generator of $\Gamma$, and the $f_{i}$ for $i=1,\ldots,k$ are polynomials of degree $\lambda_{i}$, say, which are distinguished and irreducible. Thus, $\Lambda(\Gamma)/(f_{i}^{a_{i}})$ is isomorphic to ${\mathbb{Z}}_{p}^{a_{i}\lambda_{i}}$ as abelian groups, and
@@
Furthermore, we have $\prod_{i=1}^{k}f_{i}^{a_{i}}=f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ up to a unit in $\Lambda(\Gamma)$.
We will make use of the following analogue of the semi-simplicity conjecture due to Greenberg, which was posed for elliptic curves over a number field, for abelian varieties over a global function field.
@@
Let $A$ be an abelian variety defined over a global function field $F$. Then the action of $\Gamma$ on $X(A/F^{\mathrm{cyc}})$ is completely reducible, that is, $a_{i}=1$ for all $i$ in (6.2).
We then have the following.
@@
Let $A$ be an abelian variety defined over a global function field $F$. Then we have
@@
In addition, equality holds when Conjecture 6.2 holds.
@@
We have shown in Theorem 6.1 that ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has finite kernel and cokernel. Taking the Pontryagin dual, we see that the map $X(A/F^{\mathrm{cyc}})_{\Gamma}\to X(A/F)$ has finite kernel and cokernel. It follows that
@@
Now, we have ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(X(A/F^{\mathrm{cyc}})_{\Gamma})={\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})\right)$, and since ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(\Lambda(\Gamma)/(T)\right)=1$ and $\Lambda(\Gamma)/(f,T)\simeq{\mathbb{Z}}_{p}/f(0){\mathbb{Z}}_{p}$ for a distinguished polynomial $f$, the ${\mathbb{Z}}_{p}$-rank of the latter is $1$ if and only if $T\mid f$. It thus follows from (6.2) that ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})\right)$ is equal to the number of $f_{i}$ equal to $T$. Assuming Conjecture 6.2, this is precisely to the power of $T$ dividing $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$, and the result follows. ∎

### v2

@@ 5. Akashi series
where the sum is over primes $w$ of $F^{\mathrm{cyc}}$ lying above $v$, and we also write $w$ for a fixed prime in $F_{\infty}$ lying above $w$. We claim that $J_{v}(A/F_{\infty})=0$ for $v\in S^{\prime}$. Indeed, let $w$ be a prime of $F_{\infty}$ above $v$, so that $\dim G_{w}=2$ by assumption. Then Iwasawa’s theorem ([35, Theorem 7.5.3]) implies that $F_{\infty,w}$ has no nontrivial $p$-extensions. Hence $H^{1}(F_{\infty,w},A_{p^{\infty}})=0$ and the claim follows.
Now consider $v\in S\setminus S^{\prime}$. To compute $\mathrm{Ak}\left(J_{v}(A/F_{\infty})\right)$, let $w$ be a prime of $F^{\mathrm{cyc}}$ above $v$, and fix a prime of $F_{\infty}$ above $w$, which we again denote by $w$. Let $H_{w}$ denote the decomposition subgroup of $H$ at $w$. By Shapiro’s lemma, we have
@@
Since $v\nmid p$, the inertia group at $v$ is finite, so $H_{w}$ is finite. As $G$ has no $p$-torsion, the order of $H_{w}$ is prime to $p$, and thus $\mathrm{cd}_{p}(H_{w})=0$. Therefore, the right-hand side of (5.6) vanishes for all $i\geq 1$, implying that
$H^{i}(H,J_{v}(A/F_{\infty}))=0\quad\textrm{for all $i\geq 1$.}$
For the term $J_{v}(A/F_{\infty})^{H}$, we observe that the restriction map
@@
is an isomorphism, due to the fact that $\mathrm{cd}_{p}(H_{w})=0$. In conclusion,
$\mathrm{Ak}\Big(J_{v}(A/F_{\infty})\Big)=\left\{\begin{array}[]{ll}f_{J_{v}(A/F^{\mathrm{cyc}})}&\hbox{if $v\in S\setminus S^{\prime}$,}\\ 0,&\hbox{if $v\in S^{\prime}$,}\end{array}\right$.
which implies the lemma. ∎
@@
Let $A$ be an abelian variety defined over a global function field $F$ with $\mathrm{char}(F)\neq p$, and let $F_{\infty}/F$ be an admissible $p$-adic Lie extension. Then
@@
Combining the computations from Lemmas 5.3 and 5.4 with Equation (5.3), we obtain
@@
By Theorem 4.2, we have the short exact sequence
@@
where all modules are cofinitely generated over ${\mathbb{Z}}_{p}$, by Lemma 3.3 and Theorem 3.4. Therefore,
@@
Substituting this into the earlier expression, we obtain the required conclusion of the theorem. ∎
We record an interesting corollary of our Theorem 5.5 which serves as an analogue of a result of Howson [18] (also see [7]).
@@
Retain the assumptions of Theorem 5.5. Suppose further that $G$ is pro-$p$ with dimension at least $2$. Then the following formula holds:
$\begin{split}\mathrm{rank}_{\Lambda(H)}X(A/F_{\infty})=\mathrm{rank}_{{\mathbb{Z}}_{p}}X(A/F^{\mathrm{cyc}})-\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{\mathrm{cyc}})^{\vee}\\ +\sum_{w\in S^{\prime}(F^{\mathrm{cyc}})}\mathrm{rank}_{{\mathbb{Z}}_{p}}A_{p^{\infty}}(F^{{\mathrm{cyc}}}_{w})^{\vee}.\end{split}$
@@
In the course of the proof, whenever $N$ is a finitely generated torsion $\Lambda(\Gamma)$-module with vanishing $\mu$-invariant, we shall represent the characteristic series $f_{N}$ by a distinguished polynomial. In this context, it makes sense to speak of $\deg f_{N}$. In particular, ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(N)=\deg f_{N}$. Now, suppose $M$ is a $\Lambda(G)$-module that is finitely generated over $\Lambda(H)$. Then $H_{i}(H,M)$ is finitely generated over ${\mathbb{Z}}_{p}$ for every $i$, and we have
(5.7) ${\mathrm{rank}}_{\Lambda(H)}(M)=\sum_{i\geq 0}(-1)^{i}{\mathrm{rank}}_{{\mathbb{Z}}_{p}}H_{i}(H,M)$
(cf. [18, Theorem 1.1]). Combining this with the previous observation, we see that ${\mathrm{rank}}_{\Lambda(H)}(M)$ can be computed from its Akashi series
${\mathrm{rank}}_{\Lambda(H)}(M)=\sum_{i\geq 0}(-1)^{i}\deg f_{H_{i}(H,M)}.$
Now recall the conclusion of Theorem 5.5, which we rewrite as
@@
Since $\dim H\geq 1$ by assumption, we know that ${\mathrm{rank}}_{\Lambda(H)}(A_{p^{\infty}}(F_{\infty})^{\vee})=0$. It remains to verify the identity
@@
for $v\in S^{\prime}$, where the sum runs over all primes $w$ of $F^{\mathrm{cyc}}$ above $v$. To this end, note that $J_{v}(A/F^{\mathrm{cyc}})=\oplus_{w|v}H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})$. For $v\in S^{\prime}$, the proof of Lemma 5.4 shows that $H^{1}(F^{\mathrm{cyc}}_{w},A_{p^{\infty}})=H^{1}(H_{w},A_{p^{\infty}}(F_{\infty,w}))$. Since $\dim H_{w}\geq 1$ and ${\mathrm{rank}}_{{\mathbb{Z}}_{p}[[H_{w}]]}(A_{p^{\infty}}(F_{\infty,w})^{\vee})=0$, it follows from a similar formula as (5.7) and ${\rm cd}_{p}(H_{w})=1$ that
@@ 6. Control Theorem
In this section, we will prove the following control theorem, which enables us to relate the order of vanishing at $T=0$ of the characteristic power series of ${\mathrm{Sel}}(A/F^{\mathrm{cyc}})$ to the ${\mathbb{Z}}_{p}$-corank of the Selmer group over ${\mathrm{Sel}}(A/F)$ over the base field $F$.
@@
Let $A$ be an abelian variety defined over a global function field $F$ of characteristic different from $p$. Let $F^{\mathrm{cyc}}$ denote the cyclotomic ${\mathbb{Z}}_{p}$-extension of $F$. Write $\Gamma_{n}={\mathrm{Gal}}(F^{\rm cyc}/F_{n})$, where $F_{n}$ is the $n$-th layer of the extension $F^{\mathrm{cyc}}/F$ such that $F_{n}:F]=p^{n}$. Then the restriction map
@@
has a finite kernel and cokernel, which are bounded independently of $n$.
@@
We follow closely the method used in the proof of the control theorem for fine Selmer groups over number fields, as presented in [28]. Consider the following commutative diagram with exact rows
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Sel}}(A/F_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\alpha_{n}}$${H^{1}(G_{S}(F_{n}),A_{p^{\infty}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\beta_{n}}$${\bigoplus_{v\in S}J_{v}(A/F_{n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\gamma_{n}}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${{\mathrm{Sel}}(A/F^{\rm cyc})^{\Gamma_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${H^{1}(G_{S}(F^{\rm cyc}),A_{p^{\infty}})^{\Gamma_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bigoplus_{v\in S}J_{v}(A/F^{\rm cyc})^{\Gamma_{n}}.}$
Since $\Gamma_{n}$ has $p$-cohomological dimension $1$, the inflation-restriction sequence gives:
$\ker\beta_{n}=H^{1}(\Gamma_{n},A_{p^{\infty}}(F^{\mathrm{cyc}}))\quad\textrm{and}\quad\mathrm{coker}~\!\beta_{n}=0.$
Let $M=\left(A_{p^{\infty}}(F^{\mathrm{cyc}})\right)^{\vee}$. Then $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module with a continuous $\Gamma$-action, hence a finitely generated torsion $\Lambda(\Gamma_{n})$-module for every $n$. By [35, Proposition 5.3.20],
@@
By the Mordell–Weil theorem, $M_{\Gamma_{n}}=\left(A_{p^{\infty}}(F_{n})\right)^{\vee}$ is finite, so ${\mathrm{rank}}_{{\mathbb{Z}}_{p}}M_{\Gamma_{n}}=0$. It follows that $M^{\Gamma_{n}}$ is also finite. Note that
@@
which is the Pontryagin dual of $M^{\Gamma_{n}}$. As $M$ is a finitely generated ${\mathbb{Z}}_{p}$-module, the order of $M^{\Gamma_{n}}$ is bounded by the order of $M(p)$, which is independent of $n$. It remains to show that $\ker\gamma_{n}$ is finite and bounded independently of $n$. By the snake lemma, this will imply the same uniform boundedness for $\ker\alpha_{n}$ and $\mathrm{coker}~\!\alpha_{n}$. To analyse $\gamma_{n}$, observe that $v\in S$ does not divide $p$, it suffices to show that for each prime $w$ of $F^{\mathrm{cyc}}$ lying above $v\in S$, the group $H^{1}(\Gamma_{w},A_{p^{\infty}}(F^{\mathrm{cyc}}_{w}))$ is finite and bounded independently of $n$. This can be achieved by replacing $\Gamma_{n}$ with $\Gamma_{w}$, and $M$ with $\left(A_{p^{\infty}}(F^{\mathrm{cyc}}_{w})\right)^{\vee}$, which remains a finitely generated ${\mathbb{Z}}_{p}$-module, and then applying the same argument as above. ∎
In Theorem 3.4, we showed that $X(A/F^{\mathrm{cyc}})$ is a finitely generated torsion $\Lambda(\Gamma)$-module with $\mu$-invariant zero. Thus, the structure theorem for finitely generated $\Lambda(\Gamma)$-modules implies that there exists a pseudo-isomorphism
@@
where we identify $\Lambda(\Gamma)\simeq{\mathbb{Z}}_{p}[[T]]$ using our fixed topological generator of $\Gamma$, and each $f_{i}$ for $i=1,\ldots,k$, is a distinguished irreducible polynomial of degree $\lambda_{i}$. Consequently, $\Lambda(\Gamma)/(f_{i}^{a_{i}})$ is isomorphic to ${\mathbb{Z}}_{p}^{a_{i}\lambda_{i}}$ as an abelian group, and
@@
Moreover, we have $\prod_{i=1}^{k}f_{i}^{a_{i}}=f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$ up to a unit in $\Lambda(\Gamma)$.
We will make use of the following analogue of Greenberg’s semi-simplicity conjecture ([17, Conjecture 1.12]), originally posed for elliptic curves over number fields, but here considered for abelian varieties over global function fields.
@@
Let $A$ be an abelian variety defined over a global function field $F$ with ${\rm char}(F)\neq p$. Then the action of $\Gamma$ on $X(A/F^{\mathrm{cyc}})$ is completely reducible, that is, in the decomposition (6.1), we have $a_{i}=1$ for all $i$.
We then have the following result:
@@
Let $A$ be an abelian variety defined over a global function field $F$. Then
@@
Moreover, equality holds if Conjecture 6.2 is true.
@@
By Theorem 6.1, the natural restriction map ${\mathrm{Sel}}(A/F)\to{\mathrm{Sel}}(A/F^{\mathrm{cyc}})^{\Gamma}$ has a finite kernel and cokernel, which by Pontryagin duality implies the same for the map $X(A/F^{\mathrm{cyc}})_{\Gamma}\to X(A/F)$. It follows that
@@
Next, note that
${\mathrm{rank}}_{{\mathbb{Z}}_{p}}(X(A/F^{\mathrm{cyc}})_{\Gamma})={\mathrm{rank}}_{{\mathbb{Z}}_{p}}\left(X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})\right).$
Since, for any distinguished polynomial $f(T)$, we have $\Lambda(\Gamma)/(f,T)\simeq{\mathbb{Z}}_{p}/f(0){\mathbb{Z}}_{p},$ it follows that the ${\mathbb{Z}}_{p}$-rank of $\Lambda(\Gamma)/(f,T)$ is $1$ if and only if $T\mid f$. Therefore, from (6.1), the ${\mathbb{Z}}_{p}$-rank of $X(A/F^{\mathrm{cyc}})/TX(A/F^{\mathrm{cyc}})$ is equal to the number of polynomials $f_{i}$ that are equal to $T$. Assuming Conjecture 6.2, this is precisely to the power of $T$ dividing $f_{{\mathrm{Sel}}(A/F^{\mathrm{cyc}})}$, and the result follows. ∎